I didn't know where to go for help. I took basic algebra years ago but don't really remember much. I'm trying to come up with an Excel formula that will find the value of X. I'm starting with X=(X/(Y*(Z/14)))/14. This is correct. I know I can simplify to 14X=X/(Y*(Z/14)) but then I don't know what to do with the X in the numerator. How do I get X to only one side of the equation?
Answer by jim_thompson5910(35256)

14X=X/(Y*(Z/14))

14X - X/(Y*(Z/14)) = 0

X ( 14 - 1/(Y*(Z/14)) ) = 0

X = 0 or 14 - 1/(Y*(Z/14)) = 0

Since we only care about x, the answer is X = 0

So maybe there's a typo somewhere?